Volunteer Pharmaceuticals (VP) wants to set aside money in the R&D fund to seek supplemental drug applications of its two newly developed drugs. Instead of putting away $780 million now, VP plans to set aside $A each year from year 1 through 40. What should the value of A be if VP expects to earn 7% per year for the first 5 years and 2% per year for the remaining years? Please provide formula, steps and explanation.
Answer #1

Interest rate for the first 5 years = 7%

Interest rate for the 6 to 40 years = 2%

Then,

780 = A*(P/A, 7%, 5) + A*(P/A, 2%, 35)*(P/F, 7%, 5)

780 = A*4.1002 + A*24.9986*.713

A = 780/(4.1002 + 24.9986*.713)

A = $35.58 Million

Hence, annual savings should be $35.58 Million to achieve the value of $780 Million, now.